OpenStudy (anonymous):

find the volume of a cone with radius r = 5in. and slant height l=13 in

OpenStudy (anonymous):

height of te cone by pythagoras theorem is 12. so volume=pi*r^2*h/3=3.14*25*12/3=314in^3.